The engine of a car exerts a constant pull of magnitude 500N. Find the distance traveled by the car if the work done by the force is 500kj.
1
Expert's answer
2020-04-02T10:49:54-0400

The work done by force

"W=Fd."

Hence, the distance traveled by the car

"d=\frac{W}{F}=\frac{500\times 10^3\:\rm J}{500 \:\rm N}=1000\:\rm m."